OpenStudy (anonymous):

a water park charges 6$ for admission and 1.50$ for each ride ticket. you have 36$. how many ride tickets can you buy. graph your solution.

OpenStudy (anonymous):

6 + 1.50x = 36 1.50x = 36 - 6 1.50x = 30 x = 30/1.50 x = 20 you can buy 20 tickets.